a- Rosa acquired 7 put option contracts on Imports United stock. The strike price was $42.50 and the option premium was $1.19. At expiration, Imports stock was selling for $48.93. What is the payoff on the option contracts?

b- You purchased 4 call options with a $25 strike price at a cost of $0.56 per share. On the expiration date, the underlying stock was priced at $29.97. What is the percentage return on your investment?

Answers should be formatted as a percent with 2 decimal places (e.g. 99.99).
Answer #1

a) Rosa has acquired put options, therefore she has a right to sell

Since at expiration, the stock price was $48.93, and the exercise being 42.50, she will prefer to sell at market rather than exercising the option. Therefore the option will lapse.

Benefit From Option = 0

Payoff = Benefit - Premium

Therefore there is a negative payoff of $ 8.33

b) Since Excercise price is lower than the strike price, the call option will be exercised

Benefit from option = (29.97 - 25) x 4 = 19.88

Premium Paid = 0.56 x 4 = 2.24

Therefore Payoff = 19.88 - 2.24 = 17.64